OpenStudy (anonymous):

A. write the expression 81^-3/4 in radical notation. B. evaluate the radical expression

OpenStudy (anonymous):

\[\sqrt{ 81^-3/4}= ?\] Right ?

OpenStudy (anonymous):

We can say : 81=3^4=>81^-3=(3^4)^-3 = 3^-12 and : \[\sqrt{3^-12 }=3^-12/2=3^-6\] And 4=2^2 So : \[\sqrt{2^2}=2^2/2=2\] @sexybadass22

OpenStudy (anonymous):

We have : 3^-6/2 =>3^-6=1/3^6 =>1/3^6/2=>1/3^6/2/1=1/3^6 * 2

OpenStudy (anonymous):

@sexybadass22

OpenStudy (lukecrayonz):

81^-3/4 is not sqrt(81^-3/4)

Directrix (directrix):

@CF7 is offline, I think.

OpenStudy (anonymous):

Yes .! I think this user answer is right .;)

OpenStudy (skullpatrol):